# Zhang Q Y, Xu L, Huang J H, et al

### v1

@@ 1 Introduction
Driven by the space-air-ground integrated network (SAGIN), satellite Internet of Things (S-IoT) and satellite-integrated Internet initiatives, the satellite communication (SatCom) system has witnessed a remarkable evolution, transitioning from a standalone large geostationary satellite serving mode to ultra-dense mega-constellation networks such as Starlink, Xingyun and OneWeb, catering to the vision of seamless global coverage and ubiquitous connectivity in the forthcoming sixth generation (6G) [1]. Nevertheless, as the demand for diverse intelligent applications under 6G integration continues to rise, the issue of information barriers inherent in existing independent satellite networks has become increasingly pronounced, rendering these networks incapable of providing globally open, cohesive, and unified information services. This has aroused widespread concern in academia and industry on the distributed satellite information networks (DSIN) represented by the cohesive clustered satellites (CCS) system [2, 3].
1.1 The Limitations of Current SatCom Network
To ensure interoperability and efficient utilization of SatCom, various key technologies such as advanced multi-satellite multi-beam collaborations, software defined networking (SDN), on-orbit autonomy, on-orbit computing and intelligent resource management have been proposed to achieve the goal of 6G-enabled DSIN. In addition, the standardization efforts for SatCom and Non-Terrestrial Networks (NTN) are gaining momentum, with the active participation of several standardization organizations, including the European Telecommunications Standards Institute (ETSI) and the $3^{rd}$ Generation Partnership Project (3GPP) [4]. However, current SatCom networks will not meet all the requirements of the future 6G-enabled DSIN. One of the main distinguishing features of the current SatCom systems is the prevalence of using a single satellite to provide emergency communication or Earth Observations (EO) services, which often struggle to offer continuous and robust communication services due to their singular nature and susceptibility to regional disruptions. Further, the proliferation of diverse satellite systems, including those dedicated to communication, navigation, and remote sensing, has led to a siloed approach where various departments often construct independent ground stations (GSs) to serve their specific domain applications. This fragmented infrastructure hinders the timely sharing and integrated utilization of massive satellite network business information across different satellite systems. In response, the future of the satellite networking paradigm is pivoting towards the multi-satellite collaborative shifts, i.e., the CCS system, aiming to create a cohesive platform for integrated acquisition, processing, storage, transmission, and distribution of space-based information. Last but not least, the integration of communication, sensing, and computing capabilities for distributed homogeneous or heterogeneous payloads is becoming increasingly critical. The existing satellite networks are grappling with the challenge of limited orbital and spectral resources, which are unable to scale in line with the growing complexity of space missions and the burgeoning demands of 6G-enabled DSIN. The weak functional synergy among diverse payloads, limited perception and computing capabilities, and poor dynamic coordination of heterogeneous resources impede the network from providing task/goal-oriented, information-centric and automation-level intelligent services, which is beyond what current satellite networks can deliver.
@@ 1.1 The Limitations of Current SatCom Network
The above limitations have spurred the development of DSIN, especially underscored by the maturation of micro and nanosatellite manufacturing, along with advancements in multi-satellite common orbit control, high-speed inter-satellite communication, and multi-load collaboration technologies. The typical architecture of DSIN encompasses: 1) the satellite constellations of hundreds to thousands of homogeneous satellites lacking inter-satellite control mechanisms, achieving near-real-time large-scale connectivity and seamless integration; 2) the CCS system of dozens to pairs of homogeneous or heterogeneous satellites in close proximity to form a virtual satellite, built upon orbital control, self-organizing networking and payload synergy technologies, tailored towards various complex space missions to provide flexible, reconfigurable, and resilient space-based information services [2, 3]. These innovative DSIN architectures has significantly enhanced the self-organization and self-healing capabilities of current satellite networks, which have been advanced in notable research projects around the world, i.e., NASA’s Earth Observing-1, SpaceX’s Starlink, ONION funded by European Union’s Horizon 2020, and the Fast, Flexible, Fractionated, Free-Flying (F6) launched by the Defense Advanced Research Projects Agency (DARPA) etc [5]. In this regard, the development of DSIN is no longer a mere concept but an inevitable trend in the evolution of space information technology and 6G networks.
@@ 1.2 Motivations and Contributions
2) New Air Interface and Transmission Technologies are Essential for Achieving High Spectrum Efficiency and Energy Efficiency. This includes advancements in channel modeling and estimations, cloud-native distributed MIMO cooperation and signal processing, new waveforms, multiple access approaches, multicasting mechanisms, channel coding schemes, phased array antennas, erasure transmission, high-speed inter-satellite communication, network routing, and the proper combination of all these diversity techniques. Further, another key area of focus is the development of cross-layer optimization techniques that leverage the dynamic nature of satellite networks, the characteristics of DSIN, and the diverse needs of various intelligent services. These techniques including mobility management, resource management, secure communications and testbeds, are crucial for guaranteeing high-deterministic, energy-efficient, ultra-secure and verifiable on-demand services.
@@ 2.1.1 From Transparent to Controllable/Regenerative Payload Architecture
Architecture-1: From transparent to controllable satellite: In the legacy satellite networks, given the restriction of the satellite platform, the transparent payload for communication is widely deployed, especially for the GEO and MEO systems with following restrictions, e.g.,: 1) Limited processing capability on satellites, which is usually dedicated for satellite control instead of communication payload; 2) The controlling of RF components, e.g., beam and frequency reuse mode, is limited and via the dedicated feeder link along with other information, e.g., satellite movement/gesture adjustment. Considering the hardware restrictions of satellites, the regenerative satellite payload may have partial or full BS function onboard during its evolution. Although the satellites with higher capability or lower orbits are popular for the construction of the new satellite networks, re-framing the legacy on-orbit system is still essential considering the additional aspects, e.g., cost and capability of satellite manufacture.
As shown in Figure 2, the concept of network controllable repeater (NCR) [6] is introduced to enable the evolution of legacy systems with limited processing capability onboard. For example, the legacy transparent satellite (i.e., NCR-Fwd), still only implements frequency conversion, a Radio Frequency (RF) amplifier and beam hopping in both up and down links. However, additional controlling information will be received by the NCR-MT via the control link from ground BS. In this way, the coordination direction of the communication system can be well achieved via the centralized control center to optimize the network performance according to the dynamic traffic and scheduling per UEs/Beam footprint. Moreover, this architecture is more friendly to enable the joint deployment between SatCom and cellular system, e.g., frequency sharing via dynamic turn-on/off the beam or updating the frequency used for each beam.
@@
As the exemplified network shown in Figure 3, with more capable satellites, the system can enable additional applications. For example, with the ISL, the distributed satellites in DSIN can be deployed to enable the complicated coordination onboard. For instance, enabling certain satellites to function as computation nodes for processing information, such as images, is expected to further reduce latency and alleviate traffic loads during information exchange between satellites and gateways.
@@
Additionally, in [7], a group of satellites work together to provide services to the UEs with high reliability since multiple satellites in CCS system can provide the service for the same area, along with backup in case of a satellite failure. Furthermore, several small and lightweight satellites in CCS system, e.g., CubeSats, equipped with a commercial low gain patch antenna are tuned to create a large equivalent aperture providing a huge gain and a narrow beam. The CCS system can be arranged in a free-flying configuration (i.e., wireless connected) or tethered configuration (i.e., optically connected). Simulation results presented in [8] demonstrate that the performance is comparable with the classical and distributed implementations of antenna arrays.
@@ 2.1.2 CU-DU Split Network Architecture
The ongoing work on 3GPP Rel-19 marks a milestone in the integration of satellite technologies into 5G networks as defined by 3GPP [9], particularly with the inclusion of NTN that feature regenerative, or packet-processing payloads. The regenerative satellite-based payloads offer greater flexibility and performance by hosting partial or complete functionality of a gNB, an NR BS, and supporting the Xn inter-gNB interface for ISL [10]. Nevertheless, due to the limitations in payload, power supply, and heat dissipation on satellite platforms, the processing capability of a single satellite is limited [11]. To deal with this issue, it is crucial to research and develop a distributed, open, and intelligent DSIN architecture that can provide more efficient, flexible, and smart solutions for a variety of applications, including worldwide communication, electromagnetic spectrum detection, remote sensing applications, and a multitude of other critical missions. As a promising option, the integration of open radio access networks (O-RAN) with regenerative payloads aligns with the distributed architecture that separates the Central Unit (CU) and Distributed Unit (DU) [12]. As a result, a cutting-edge CU-DU Split [13] network architecture can be implemented. It originates from terrestrial 5G networks, where the CU is tasked with managing non-real-time protocols and services, including Radio Resource Control (RRC) and Packet Data Convergence Protocol (PDCP). Typically situated in data centers or regional hubs, the CU facilitates efficient resource management and coordination. Meanwhile, the DU is charged with the real-time processing of wireless access layer protocols, such as Medium Access Control (MAC) and Radio Link Control (RLC). It is connected to the Radio Unit (RU), often located in close proximity or co-located with the RU to minimize latency. The CU-DU Split not only fosters the sharing of baseband resources but also paves the way for the slicing and cloudification of wireless access. This innovative approach ensures effective collaboration between sites in complex DSIN. In addition, this separation is pivotal for unlocking the full potential of 6G networks in space, where the NTN gateway at the end of the feeder link can function as a router to the core network. The ISLs are indispensable for regenerative payloads, particularly in Non-Geostationary Satellite Orbits (NGSO) such as LEO and MEO, where the satellites move significantly faster than the Earth’s rotation. This rapid movement necessitates seamless handovers of the feeder link to alternative gateways and the maintenance of service continuity by different space-borne platforms.
@@
For satellite systems, the Satellite Centralized Unit (SAT-CU) serves as the centralized processing unit, enabling efficient sharing of processing capabilities. By centralizing resource management and complex data processing in the SAT-CU, the system can dynamically allocate resources and quickly adjust task configurations to meet changing demands, while coordinating the collaborative work of multiple Satellite Distributed Units (SAT-DU) in CCS system to enhance the overall efficiency of DSIN. The SAT-CU, in its role as the control plane and user data anchor, can reduce data interruption delays caused by handovers, further improving the user experience. Meanwhile, the SAT-DU is responsible for preliminary data processing and filtering with high real-time requirements, reducing data transmission delays and improving response speed. Since the SAT-DU only needs to provide RF and limited baseband processing capabilities, with user data being collected and processed centrally by SAT-CU without the need for independent core network connections, the weight, power consumption, and complexity of SAT-DU satellites can be effectively reduced, thereby further reducing the comprehensive cost of satellite network construction and deployment. This architecture also improves the robustness and reliability of the CCS system, better isolates and handles faults, and reduces the risk of single-point failures.
@@ 2.2 Distributed Satellite Computing Network Architecture
With the development of satellite technology, the available computing, communication, sensing and storage resources of satellites are increasingly abundant. To reduce the data transmitted back to the ground for processing and improve information delivery efficiency, particularly in tasks such as pre-processing and target recognition of remote sensing images, the requirements of intensive computation tasks executed at satellites have become more urgent. The consensus among industry and academics is that by leveraging the regenerative capabilities of satellites, deploying computational power nodes on satellites represents a promising solution [17]. However, how to efficiently schedule the heterogeneous resources at distributed satellites and make those satellites located in a wide area collaborative is still an urgent problem that needs to be solved.
@@ 2.2.1 Key Technologies of Satellite Computing Deployment
Figure 5: Network Architecture of Integrated Sensing, Communication, Computation, Control, and Storage for DSIN
@@ 2.3 Reconfigurable Satellite Formation Flying
As a key advancement in DSIN, satellite formation flying (SFF) plays a pivotal role in modern space exploration and applications to maintain a targeted orbit configuration with desired relative separation and orientation between multiple spacecraft. The concept of SFF was first introduced in the 1970s to leverage multiple satellites to conduct interferometric infrared synthetic aperture imaging tasks traditionally handled by single, large satellites [27]. A specific satellite formation is composed of multiple satellites distributed on the same or adjacent orbits, functioning as an extended “virtual spacecraft” and avoiding the technical and financial challenges of building one satellite of equivalent size. For a specific SFF, the member satellites are relatively close to each other, with fixed relationships or dynamical relationships that can be represented by linearized equations, and the space-relative positions are commonly expressed using Cartesian coordinate parameters. Early satellite formations are primarily employed in Synthetic Aperture Radar (SAR) applications, such as the U.S. Techsat-21 program, France’s Cartwheel program, and Germany’s Pendulum project [28]. Due to their configuration flexibility, satellite formations are widely utilized in Earth environment monitoring missions, such as the ESA Swarm, QB-50, and GRACE projects. In space science experiments, many projects such as TPF and LiSA have been launched, and satellite formations for gravitational wave detection projects like Tianqin and Taiji are under construction in China.
@@
The SFF evolved from simple dual-satellite configurations to more complex architectures involving dozens or even hundreds of small satellites. A satellite swarm (or cluster) is a broad concept. In general, any satellite system in which multiple satellites collaborate to perform tasks can be referred to as a satellite cluster. To distinguish from other concepts, the “satellite swarm” discussed in this subsection typically consists of multiple satellites with similar orbital parameters and close spatial proximity [27]. The concept of “satellite clusters” emphasizes extending system functionality on a spatial scale, without requiring the satellites to maintain stable relative positions. The functionality of a satellite swarm can be realized by ensuring a certain number of satellites within a specified spatial range. In remote sensing applications, projects like OLFAR, SWIFT, and KickSat propose the deployment of large numbers of satellites as observation arrays to improve remote sensing performance. Remote sensing satellite swarms such as FISC and Flock enable frequent global observation updates.
@@
A “satellite constellation” is a satellite system composed of multiple satellites distributed in several identically shaped orbits according to a specific pattern. Generally, the satellites in a constellation are evenly distributed along their orbits, with nearly equal spacing between orbital planes [28]. The concept of satellite constellations emphasizes a balanced spatial distribution of satellite functionality. The relative positions of satellites within a constellation are typically characterized by orbit parameters, and the use of homologous orbits ensures a stable configuration, thereby reducing the need for frequent maintenance. Satellite constellations are commonly employed for global or latitudinally distributed communication and navigation services. Since the 1970s, various countries have progressively developed mid- and high-orbit navigation constellations such as GPS, BeiDou, GLONASS, and Galileo, low-orbit communication constellations such as Iridium and Global Star, as well as inclined elliptical orbit constellations such as Molniya and O3B. Large-scale LEO Internet constellations, represented by Starlink, OneWeb, Kuiper, and Telesat, have also been extensively deployed. The aforementioned satellite systems all use fully functional satellites as their basic units. Additionally, there is a distribution concept where satellite subsystems are separated and perform formation flying in space.
@@
To meet the demands of different scenarios, the CCS systems require scalable and reconfigurable capabilities for configuration adjustments. In near-circular orbit formation flying scenarios, the relative motion of satellites can be described by the linearized Hill equation (also known as the C-W equation) [29]. Building on this, Tschauner and Hempel [30], through variable substitution, proposed the T-H equation, which is a linear time-varying equation describing elliptical orbits via the eccentric Anomaly. The aforementioned linearized equations are valid only for formation scenarios where the inter-satellite distance is less than 20 km. However, in non-near-circular orbits and conservative perturbation environments, obtaining the analytical solutions of these equations is challenging, rendering them unsuitable for relative configuration design, but more suitable for motion control when combined with closed-loop control algorithms such as PID, sliding mode, optimization, or robust control methods. In addition, driven by the growing need for large-scale, the DSIN is envisioned to provide continuous, real-time services from deep space exploration to planetary orbit application fields [31], such as the high-orbit high-resolution optical EO, space target attachment, electromagnetic force formation and Confederacy space system design, etc. As such, the development of satellite formation configurations has evolved from simple relative position maintenance to complex multi-satellite cooperation, and further to intelligent and adaptive control. However, these expanded formations introduced new challenges in terms of communication constraints, incomplete information, and the need for advanced coordination strategies to maintain operational efficiency.
@@
Various approaches to satellite formation control have been explored in the literature, each offering unique advantages and addressing specific challenges associated with maintaining and controlling satellite configurations in space, as shown in Figure 6. One of the foundational approaches to formation control is the leader-following method. In this configuration, one or more satellites serve as “leaders,” setting the trajectory for the rest of the satellites, or “followers,” to track [32]. The followers adjust their position and orientation relative to the leader, thus achieving a coordinated formation. The Leader-following methods can be implemented in various structures, including single-leader, multi-leader, and virtual-leader formations, which offer different levels of flexibility and robustness. However, a key limitation of the leader-following approach is its dependence on the leader. If the leader experiences a malfunction or significant perturbations, the entire formation may become destabilized, posing risks to the mission [33].
@@
In constraint to centralized formation control, the behavior-based formation control method represents a more decentralized approach [34]. Instead of relying on a designated leader, this method assigns specific behaviors to each satellite, such as collision avoidance, formation keeping, and reconfiguration. Through these individual behaviors and local control rules, the overall formation is maintained in a coordinated manner. The core challenge in behavior-based control lies in designing effective behavior coordination mechanisms that ensure global formation objectives are achieved. For example, satellites may have to balance conflicting requirements, such as maintaining formation while avoiding potential collisions [35]. Behavior-based methods are particularly advantageous for large formations with multiple interacting units, as they allow for greater adaptability and autonomy. However, the decentralized nature of this approach can make it difficult to achieve the high levels of precision required in certain missions, as the formation’s overall stability depends on the interaction of individual satellite behaviors.

### v2

@@ 1 Introduction
Driven by the space-air-ground integrated network (SAGIN), satellite Internet of Things (S-IoT) and satellite-integrated Internet initiatives, the satellite communication (SatCom) system has witnessed a remarkable evolution, transitioning from a standalone large geostationary satellite serving mode to ultra-dense mega-constellation networks such as Starlink, Xingyun and OneWeb, catering to the vision of seamless global coverage and ubiquitous connectivity in the forthcoming sixth generation (6G) [1]. Nevertheless, as the demand for diverse intelligent applications in 6G integration continues to rise, the issue of information barriers inherent in existing independent satellite networks has become increasingly pronounced, rendering these networks incapable of providing globally open, cohesive, and unified information services. This has raised widespread concern in academia and industry regarding the distributed satellite information networks (DSIN) represented by the cohesive clustered satellites (CCS) system [2, 3].
1.1 The Limitations of Current Satellite Network
[…]
@@ 1.1 The Limitations of Current Satellite Network
The above limitations have spurred the development of DSIN, especially highlighted by the maturation of micro and nanosatellite manufacturing, along with advancements in multi-satellite common orbit control, high-speed inter-satellite communication, and multi-load collaboration technologies. The typical architecture of DSIN encompasses: 1) the satellite constellations of hundreds to thousands of homogeneous satellites lacking inter-satellite control mechanisms, achieving near-real-time large-scale connectivity and seamless integration; 2) the CCS system of dozens to pairs of homogeneous or heterogeneous satellites in close proximity to form a virtual satellite, built upon orbital control, self-organizing networking and payload synergy technologies, tailored towards various complex space missions to provide flexible, reconfigurable, and resilient space-based information services [2, 3]. These innovative DSIN architectures has significantly enhanced the self-organization and self-healing capabilities of current satellite networks, which have been advanced in notable research projects around the world, i.e., NASA’s Earth Observing-1, SpaceX’s Starlink, ONION funded by European Union’s Horizon 2020, and the Fast, Flexible, Fractionated, Free-Flying (F6) launched by the Defense Advanced Research Projects Agency (DARPA) etc [5]. In this regard, the development of DSIN is no longer a mere concept but an inevitable trend in the evolution of space information technology and 6G networks.
@@ 1.2 Motivations and Contributions
2) New Air Interface and Transmission Technologies are Essential for Achieving High Spectrum Efficiency and Energy Efficiency. This includes advancements in channel modeling and estimation, cloud-native distributed MIMO cooperation and signal processing, new waveforms, multiple access approaches, multicasting mechanisms, channel coding schemes, phased array antennas, erasure transmission, high-speed inter-satellite communication, network routing, and the proper combination of all these diversity techniques. Further, another key area of focus is the development of cross-layer optimization techniques that leverage the dynamic nature of satellite networks, the characteristics of DSIN, and the diverse needs of various intelligent services. These techniques, including mobility management, resource management, secure communications and testbeds, are crucial for guaranteeing high-deterministic, energy-efficient, ultra-secure and verifiable on-demand services.
@@ 2.1.1 From Transparent to Controllable/Regenerative Payload Architecture
Architecture-1: From transparent to controllable satellite: In legacy satellite networks, given the restriction of the satellite platform, the transparent payload for communication is widely deployed, especially for the GEO and MEO systems with following restrictions, e.g.,: 1) Limited processing capability on satellites, which is usually dedicated for satellite control instead of communication payload; 2) The controlling of RF components, e.g., beam and frequency reuse mode, is limited and via the dedicated feeder link along with other information, e.g., satellite movement/gesture adjustment. Considering the hardware restrictions of satellites, the regenerative satellite payload may have partial or full BS function onboard during its evolution. Although the satellites with higher capability or lower orbits are popular for the construction of the new satellite networks, re-framing the legacy on-orbit system is still essential considering the additional aspects, e.g., cost and capability of satellite manufacture.
As shown in Figure 2, the concept of network controllable repeater (NCR) [6] is introduced to enable the evolution of legacy systems with limited processing capability onboard. For example, the legacy transparent satellite (i.e., NCR-Fwd), still only implements frequency conversion, a Radio Frequency (RF) amplifier, and beam hopping in both up and down links. However, additional controlling information will be received by the NCR-MT via the control link from ground BS. In this way, the coordination direction of the communication system can be well achieved via the centralized control center to optimize the network performance according to the dynamic traffic and scheduling per UEs/Beam footprint. Moreover, this architecture is more friendly to enable the joint deployment between SatCom and cellular system, e.g., frequency sharing via dynamic turn-on/off the beam or updating the frequency used for each beam.
@@
As the exemplified network shown in Figure 3, with more capable satellites, the system can enable additional applications. For example, with the ISL, the distributed satellites in DSIN can be deployed to enable complicated coordination onboard. For instance, enabling certain satellites to function as computation nodes for processing information, such as images, is expected to further reduce latency and alleviate traffic loads during information exchange between satellites and gateways.
@@
Additionally, in [7], a group of satellites work together to provide services to the UEs with high reliability since multiple satellites in CCS system can provide the service for the same area, along with backup in case of a satellite failure. Furthermore, several small and lightweight satellites in CCS system, e.g., CubeSats, equipped with a commercial low gain patch antenna are tuned to create a large equivalent aperture providing a huge gain and a narrow beam. The CCS system can be arranged in a free-flying configuration (i.e., wireless connected) or tethered configuration (i.e., optically connected). The Simulation results presented in [8] demonstrate that the performance is comparable with the classical and distributed implementations of antenna arrays.
@@ 2.1.2 CU-DU Split Network Architecture
The ongoing work on 3GPP Rel-19 marks a milestone in the integration of satellite technologies into 5G networks as defined by 3GPP [9], particularly with the inclusion of NTN that feature regenerative, or packet-processing payloads. The regenerative satellite-based payloads offer greater flexibility and performance by hosting partial or complete functionality of a gNB, an NR BS, and supporting the Xn inter-gNB interface for ISL [10]. Nevertheless, due to the limitations in payload, power supply, and heat dissipation on satellite platforms, the processing capability of a single satellite is limited [11]. To deal with this issue, it is crucial to research and develop a distributed, open and intelligent DSIN architecture that can provide more efficient, flexible and smart solutions for a variety of applications, including worldwide communication, electromagnetic spectrum detection, remote sensing applications, and a multitude of other critical missions. As a promising option, the integration of open radio access networks (O-RAN) with regenerative payloads aligns with the distributed architecture that separates the Central Unit (CU) and Distributed Unit (DU) [12]. As a result, a cutting-edge CU-DU Split [13] network architecture can be implemented. It originates from terrestrial 5G networks, where the CU is tasked with managing non-real-time protocols and services, including Radio Resource Control (RRC) and Packet Data Convergence Protocol (PDCP). Typically situated in data centers or regional hubs, the CU facilitates efficient resource management and coordination. Meanwhile, the DU is charged with the real-time processing of wireless access layer protocols, such as Medium Access Control (MAC) and Radio Link Control (RLC). It is connected to the Radio Unit (RU), often located in close proximity or co-located with the RU to minimize latency. The CU-DU Split not only fosters the sharing of baseband resources but also paves the way for the slicing and cloudification of wireless access. This innovative approach ensures effective collaboration between sites in complex DSIN. In addition, this separation is pivotal for unlocking the full potential of 6G networks in space, where the NTN gateway at the end of the feeder link can function as a router to the core network. The ISLs are indispensable for regenerative payloads, particularly in Non-Geostationary Satellite Orbits (NGSO) such as LEO and MEO, where the satellites move significantly faster than the Earth’s rotation. This rapid movement necessitates seamless handovers of the feeder link to alternative gateways and the maintenance of service continuity by different space-borne platforms.
@@
For satellite systems, the Satellite Centralized Unit (SAT-CU) serves as the centralized processing unit, enabling efficient sharing of processing capabilities. By centralizing resource management and complex data processing in the SAT-CU, the system can dynamically allocate resources and quickly adjust task configurations to meet changing demands, while coordinating the collaborative work of multiple Satellite Distributed Units (SAT-DU) in CCS system to enhance the overall efficiency of DSIN. The SAT-CU, in its role as the control plane and user data anchor, can reduce data interruption delays caused by handovers, further improving the user experience. Meanwhile, the SAT-DU is responsible for preliminary data processing and filtering with high real-time requirements, reducing data transmission delays and improving response speed. Since SAT-DU only needs to provide RF and limited baseband processing capabilities, with user data being collected and processed centrally by SAT-CU without the need for independent core network connections, the weight, power consumption, and complexity of SAT-DU satellites can be effectively reduced, thereby further reducing the comprehensive cost of satellite network construction and deployment. This architecture also improves the robustness and reliability of the CCS system, better isolates and handles faults, and reduces the risk of single-point failures.
@@ 2.2 Distributed Satellite Computing Network Architecture
With the development of satellite technology, the available computing, communication, sensing and storage resources of satellites are increasingly abundant. To reduce the data transmitted back to the ground for processing and improve information delivery efficiency, particularly in tasks such as pre-processing and target recognition of remote sensing images, the requirements of intensive computation tasks executed at satellites have become more urgent. The consensus among industry and academics is that by leveraging the regenerative capabilities of satellites, deploying computational power nodes on satellites represents a promising solution [17]. In this context, distributed satellite computing has become a common approach for onboard large-scale and massive computation tasks, offering benefits such as high reliability, fault tolerance, scalability, and fast computation. However, distributed satellite computing faces significant challenges. On the one hand, the computing nodes in the distributed satellite network must exchange numerous intermediate results with each other to compute the final result, which greatly increases communication overhead, especially in high-dynamics, high-mobility LEO satellite networks. On the other hand, distributed satellite computing is executed by a large number of computing nodes, which may have varying computing and networking resources. As a result, there can be straggling nodes—computing nodes that run slower than others, unintentionally increasing the overall time required to complete the computing tasks. To address these challenges, the concept of coded distributed computing, which combines coding techniques and distributed computing, has been proposed by [18] and has recently garnered significant attention. By employing coding-theoretic techniques, the distributed computing framework encodes the subtasks for the computing nodes, enabling the master node to recover the final result from partially finished nodes, thus mitigating the effects of stragglers [19]. Furthermore, coded distributed computing is highly compatible with NFV, mobile edge computing [20], and collaborative deep reinforcement learning (DRL) [21], which facilitates unified resource and service management within SAGIN. The network topology in DSIN is characterised by strong dynamic properties, with asymmetric node computing capabilities, making it difficult to apply AI methods designed for terrestrial networks to DSIN. To address this, [21] proposes a collaborative DRL paradigm for satellite-terrestrial networks to achieve intelligent multi-orbit spectrum and computing resource management. This paradigm employs a DRL algorithm combining graph convolutional networks and federated learning (FL) to analyze and infer network topology features, guiding the communication and computing resource management of distributed satellites. This enables efficient adaptive computing offloading while avoiding model retraining due to node mobility or failures. Additionally, the integration of data centers into satellite networks has also attracted attention [3] for achieving low latency, flexible networking, and efficient resource utilisation. Moreover, by deeply integrating cloud computing, the cloud-native satellite computing network can shape DSIN into a distributed, elastic, and horizontally scalable satellite system composed of interrelated microservices, isolating state in a minimal number of stateful components [22]. However, how to efficiently schedule the heterogeneous resources at distributed satellites and make those satellites located in a wide area collaborative is still an urgent problem that needs to be solved.
@@ 2.2.1 Key Technologies of Satellite Computing Deployment
Figure 5: Distributed satellite computing network architecture for DSIN.
@@ 2.3 Reconfigurable Satellite Formation Flying
As a key advancement in DSIN, satellite formation flying (SFF) plays a pivotal role in modern space exploration and applications to maintain a targeted orbit configuration with desired relative separation and orientation between multiple spacecraft. The concept of SFF was first introduced in the 1970s to leverage multiple satellites to conduct interferometric infrared synthetic aperture imaging tasks traditionally handled by single, large satellites [32]. A specific satellite formation is composed of multiple satellites distributed on the same or adjacent orbits, functioning as an extended “virtual spacecraft” and avoiding the technical and financial challenges of building one satellite of equivalent size. For a specific SFF, the member satellites are relatively close to each other, with fixed relationships or dynamical relationships that can be represented by linearized equations, and the space-relative positions are commonly expressed using Cartesian coordinate parameters. Early satellite formations are primarily employed in Synthetic Aperture Radar (SAR) applications, such as the US Techsat-21 program, France’s Cartwheel program, and Germany’s Pendulum project [33]. Due to their configuration flexibility, satellite formations are widely utilized in Earth environment monitoring missions, such as the ESA Swarm, QB-50, and GRACE projects. In space science experiments, many projects such as TPF and LiSA have been launched, and satellite formations for gravitational wave detection projects like Tianqin and Taiji are under construction in China.
@@
The SFF evolved from simple dual-satellite configurations to more complex architectures involving dozens or even hundreds of small satellites. A satellite swarm (or cluster) is a broad concept. In general, any satellite system in which multiple satellites collaborate to perform tasks can be referred to as a satellite cluster. To distinguish from other concepts, the “satellite swarm” discussed in this subsection typically consists of multiple satellites with similar orbital parameters and close spatial proximity [32]. The concept of “satellite clusters” emphasizes extending the functionality of system on a spatial scale, without requiring the satellites to maintain stable relative positions. The functionality of a satellite swarm can be realized by ensuring a certain number of satellites within a specified spatial range. In remote sensing applications, projects like OLFAR, SWIFT, and KickSat propose the deployment of large numbers of satellites as observation arrays to improve remote sensing performance. Remote sensing satellite swarms such as FISC and Flock enable frequent global observation updates.
@@
A “satellite constellation” is a satellite system composed of multiple satellites distributed in several identically shaped orbits according to a specific pattern. Generally, satellites in a constellation are evenly distributed along their orbits, with nearly equal spacing between orbital planes [33]. The concept of satellite constellations emphasizes a balanced spatial distribution of satellite functionality. The relative positions of satellites within a constellation are typically characterized by orbit parameters, and the use of homologous orbits ensures a stable configuration, thereby reducing the need for frequent maintenance. Satellite constellations are commonly employed for global or latitudinally distributed communication and navigation services. Since the 1970s, various countries have progressively developed mid- and high-orbit navigation constellations such as GPS, BeiDou, GLONASS, and Galileo, low-orbit communication constellations such as Iridium and Global Star, as well as inclined elliptical orbit constellations such as Molniya and O3B. Large-scale LEO Internet constellations, represented by Starlink, OneWeb, Kuiper, and Telesat, have also been extensively deployed. The aforementioned satellite systems all use fully functional satellites as their basic units. Additionally, there is a distribution concept where satellite subsystems are separated and perform formation flying in space.
@@
To meet the demands of different scenarios, CCS systems require scalable and reconfigurable capabilities for configuration adjustments. In near-circular orbit formation flying scenarios, the relative motion of satellites can be described by the linearized Hill equation (also known as the C-W equation) [34]. Building on this, [35] proposed the T-H equation through variable substitution, which is a linear time-varying equation describing elliptical orbits via the eccentric Anomaly. The aforementioned linearized equations are valid only for formation scenarios where the inter-satellite distance is less than 20 km. However, in non-near-circular orbits and conservative perturbation environments, obtaining the analytical solutions of these equations is challenging, rendering them unsuitable for relative configuration design or the construction of edge information hubs [36], but more suitable for motion control when combined with closed-loop control algorithms such as PID, sliding mode, optimization, or robust control methods. In addition, driven by the growing need for large-scale, the DSIN is envisioned to provide continuous, real-time services from deep space exploration to planetary orbit application fields [37], such as the high-orbit high-resolution optical EO, space target attachment, electromagnetic force formation and Confederacy space system design, etc. As such, the development of satellite formation configurations has evolved from simple relative position maintenance to complex multi-satellite cooperation, and further to intelligent and adaptive control. However, these expanded formations introduced new challenges in terms of communication constraints, incomplete information, and the need for advanced coordination strategies to maintain operational efficiency.
@@
Various approaches to satellite formation control have been explored in the literature, each offering unique advantages and addressing specific challenges associated with maintaining and controlling satellite configurations in space, as shown in Figure 6. One of the foundational approaches to formation control is the leader-follower method. In this configuration, one or more satellites serve as “leaders,” setting the trajectory for the rest of the satellites, or “followers,” to track [38]. The followers adjust their position and orientation relative to the leader, thus achieving a coordinated formation. The Leader-following methods can be implemented in various structures, including single-leader, multi-leader, and virtual-leader formations, which offer different levels of flexibility and robustness. However, a key limitation of the leader-following approach is its dependence on the leader. If the leader experiences a malfunction or significant perturbations, the entire formation may become destabilized, posing risks to the mission [39].
@@
In constraint to centralized formation control, the behavior-based formation control method represents a more decentralized approach [40]. Instead of relying on a designated leader, this method assigns specific behaviors to each satellite, such as collision avoidance, formation keeping, and reconfiguration. Through these individual behaviors and local control rules, the overall formation is maintained in a coordinated manner. The core challenge in behavior-based control lies in designing effective behavior-coordination mechanisms that ensure global formation objectives are achieved. For example, satellites may have to balance conflicting requirements, such as maintaining formation while avoiding potential collisions [41]. Behavior-based methods are particularly advantageous for large formations with multiple interacting units, as they allow for greater adaptability and autonomy. However, the decentralized nature of this approach can make it difficult to achieve the high levels of precision required in certain missions, as the formation’s overall stability depends on the interaction of individual satellite behaviors.